# CROSS-GAiT: Cross-Attention-Based Multimodal Representation Fusion for Parametric Gait Adaptation in Complex Terrains

### v1

@@ IV-D Model for Gait Parameter Generation
We pass the combined latent representation $\mathbf{Z}^{t}_{\text{comb}}$ through a multi-layer perceptron (MLP) regressor to predict suitable gait parameters ($\mathbf{g}$), specifically, the robot’s step height and hip splay. We train the regressor using the MSE loss, based on the training labels generated using the analysis of (5) on the dataset.
@@ V-A Implementation
We implemented CROSS-GAiT using PyTorch and trained all models on an Nvidia A5000 GPU. For real-world experiments, we used the Ghost Vision 60 robot from Ghost Robotics, equipped with a front-facing wide-angle camera and an onboard Intel NUC 11 (Intel i7 CPU, NVIDIA RTX 2060 GPU). Our system runs in real-time at 60Hz. Path planning was implemented using [4]. We sample a 6-channel IMU and joint effort data, comprising 12 channels (one for each motor, measured in Nm), at 25 Hz. We combine these data streams to form the input time series data $\mathbf{X}^{t}_{\text{TS}}$, resulting in $n_{\text{ch}}=18$. $d_{\text{TS}}=100$, is chosen empirically to capture the frequency components of the robot’s cyclic gaits. $d_{m}=160$ based on [46].
@@ V-E Analysis and Discussion
Comparison on Embedding space: In Table I, we compare the terrain classification performance of the embeddings generated by CROSS-GAiT and STERLING [6]. CROSS-GAiT achieves higher accuracy, precision, recall, and F1-score. By combining visual and time-series data through cross-attention, CROSS-GAiT generates richer latent representations, enhancing the classifier’s ability to distinguish between diverse terrains.

### v2

@@ IV-D Model for Gait Parameter Generation
We pass the combined latent representation $\mathbf{Z}^{t}_{\text{comb}}$ through a multi-layer perceptron (MLP) regressor to predict suitable gait parameters ($\mathbf{g}$), specifically, the robot’s step height and hip splay. We train the regressor using the MSE loss, based on the training labels generated using the analysis of (5) on the dataset. During this phase, all layers preceding the regressor, including the Masked Autoencoder, Dilated Causal Convolutional Encoder, and cross-attention/sel-attention layers, are kept frozen.
@@ V-A Implementation
We implemented CROSS-GAiT using PyTorch and trained all models on an Nvidia A5000 GPU. For real-world experiments, we used the Ghost Vision 60 robot from Ghost Robotics, equipped with a front-facing wide-angle camera and an onboard Intel NUC 11 (Intel i7 CPU, NVIDIA RTX 2060 GPU). Our system operates in real-time at 60Hz on this hardware configuration. Path planning was implemented using [4]. We sample a 6-channel IMU and joint effort data, comprising 12 channels (one for each motor, measured in Nm), at 25 Hz. We combine these data streams to form the input time series data $\mathbf{X}^{t}_{\text{TS}}$, resulting in $n_{\text{ch}}=18$. $d_{\text{TS}}=100$, is chosen empirically to capture the frequency components of the robot’s cyclic gaits. $d_{m}=160$ based on [46].
@@ V-E Analysis and Discussion
Comparison on Embedding space: In Table I, we compare the terrain classification performance of the embeddings generated by CROSS-GAiT and STERLING [6] using a support vector classifier. CROSS-GAiT achieves higher accuracy, precision, recall, and F1-score. By combining visual and time-series data through cross-attention, CROSS-GAiT generates richer latent representations, enhancing the classifier’s ability to distinguish between diverse terrains.
@@ VI Conclusions, Limitations, and Future Work
Appendix A Appendix
A-A Hyperparameters Used During Training
The training process for the CROSS-GAiT model involves multiple stages, where different components are trained separately and then combined. The common hyperparameters for each component in CROSS-GAiT are summarized table III. Furthermore, the specific hyperparameters for each method are summarized in table IV.
Hyperparameter Masked Autoencoder Dilated Causal Conv. Encoder Cross-Attention Fusion Fully Connected Network
Learning Rate 0.00025 0.0001 0.00005 0.0001
Batch Size 128 256 1024 1024
Optimizer AdamW Adam AdamW AdamW
Epochs 400 10 100 5
Table III: Common hyperparameters across different components of the CROSS-GAiT model.
Masked Autoencoder
Hyperparameter Value Embedding Dimension 768 Masking Ratio 0.75 Weight Decay 0.05
Dilated Causal Conv. Encoder
Hyperparameter Value Dilation Rate 2 Kernel Size 3
Cross-Attention
Hyperparameter Value Embedding Dimension 160 Dropout 0.1 Weight Decay 0.01 Sup Con tempereature 0.05
Table IV: Individual hyperparameters for each component of the CROSS-GAiT model.
A-B Dynamic Window for Gait Adaptation
To ensure smooth transitions in gait parameters (step height and hip splay) while maintaining control over the robot’s movements, a dynamic window mechanism is implemented. This mechanism limits the changes in step height and hip splay per iteration, as well as enforcing maximum and minimum values for both parameters to avoid motor control issues caused by hardware limitations.
A-B1 Maximum and Minimum Values
The system enforces both maximum and minimum allowable values for step height and hip splay to avoid exceeding motor capabilities, ensuring safe and controllable movement.
  • Maximum Step Height: 0.3 meters
  • Minimum Step Height: 0.03 meters
  • Maximum Hip Splay: 0.2 meters
  • Minimum Hip Splay: 0.05 meters
A-B2 Maximum Change Per Iteration
The dynamic window also restricts the change in both step height and hip splay per iteration, allowing the system to adapt gradually without causing instability due to sudden shifts in gait.
  • Maximum Step Height Change Per Iteration: 0.01 meters
  • Maximum Hip Splay Change Per Iteration: 0.01 meters
A-C Data Collection Processs (extending Section V-B)
The data was collected under various lighting conditions, such as bright sunlight, overcast skies, and shaded areas. These varying conditions were crucial for ensuring that the visual model could generalize across different environments and remain robust when encountering challenges like shadows or direct sunlight, which could otherwise affect the perception of terrain.
Since the data collection was restricted to well-defined patches of specific terrain types (e.g., grass, hard surfaces, sand, rocks, and pebbles), there was no need to manually label individual images. The entire segment of images collected in a particular terrain zone was uniformly labeled, simplifying the process while maintaining data quality.
Additionally, to capture the full range of the robot’s dynamics, the data was recorded at various walking speeds. This approach allowed the model to learn how different gaits performed across a range of speeds, from slow and deliberate movements to faster, more dynamic gaits. These variations also enriched the IMU and joint effort data, providing a more comprehensive representation of the robot’s interaction with diverse terrains.